OpenStudy (anonymous):

PLEASE PLEASE HELP ME ASAP....An investor invested a total of 1,900 in two mutual funds. One fund earned a 4% profit while the other earned a 2% profit. If the investors total profit was $58, how much was invested in each mutal fund 9

OpenStudy (anonymous):

I dont know

OpenStudy (mertsj):

x = amount invested in 4% fund y= amount invested in 2% fund .04x = interest from 4% fund .02x = profit from 2% fund x+y=1900 .04x+.02y=58

OpenStudy (mertsj):

Solve that system

OpenStudy (anonymous):

so 6???

OpenStudy (mertsj):

Post your work so I can see how you got 6. Is that for x or y?

OpenStudy (anonymous):

i did 2+4=6.....idk i dnt get this

OpenStudy (mertsj):

x+y=1900 Solve that for x

OpenStudy (anonymous):

1000+900

OpenStudy (mertsj):